This problem involves using supply and demand equations.

Suppose the market is defined by

Demand: Q = 164 – 3P

Supply: Q = 2 + 3P

At a price of P = 17, what is the size of the shortage that will exist in the market?

Type in the amount of the shortage at price P = 17.  Once again, just enter in the number, and do not put in any other text or symbols.
Answer #1

Answer

P=$17

Qd=164-3*17

=113

Qs=2+3*17

=53 units

Shortage =Qd-Qs

=113-53

=60

the shortage at price P = 17 is 60 units.